OpenStudy (anonymous):

CAL 3 Problem Using Lagrange multipliers.

OpenStudy (anonymous):

Find the critical points of \[f(x,y,z)=x+2y-z\] on the surface \[z=1-x^2-y^2\] Check by substitution of the constraint into the function

OpenStudy (tkhunny):

Have you considered substitution? \(f(x,y) = x + 2y - (1 - x^{2} - y^{2})\)

OpenStudy (anonymous):

Sorry, I didn't put the complete question. I made the correction

OpenStudy (anonymous):

I have to solve using Lagrange multipliers

OpenStudy (tkhunny):

Fair enough. With substitution, we went down one variable, With Lagrange, we go up one. \(f(x,y,z,\lambda) = x + 2y - z + \lambda(x^{2} + y^{2} + z - 1)\) Okay, now make all four of those partial derivates zero at the same time!

OpenStudy (anonymous):

how did you find \[f(x,y,z, \lambda)\] ?

OpenStudy (tkhunny):

Your constraint is \(g(x,y,z) = c\). You must see this. \(f(x,y,z,\lambda) = f(x,y,z) + \lambda (g(x,y,z) - c)\) It's a relatively simple construction once you see it.

OpenStudy (anonymous):

Thank you for explaining, thats exactly how i thought you found it. so the four partials would be: \[f _{x}=1+2\lambda x\] \[f _{y}=2+2\lambda y\] \[f _{z}=1+\lambda\] \[f _{\lambda}= x^2+y^2+z-1\]

OpenStudy (anonymous):

four equations and four unknowns. do i then solve for lambda by fsubx - fsuby

OpenStudy (tkhunny):

You must first get \(f_{z}\) correct. ;-)

OpenStudy (anonymous):

ohh -1

OpenStudy (tkhunny):

It's kind of traditional to solve the first three for x, y, and z in terms of \(\lambda\) and substitute into the fourth.

OpenStudy (anonymous):

this is what i came up with \[x=-\frac{ 1 }{ 2\lambda }\] \[y=-\frac{ 1 }{ \lambda }\] but no "z" in the third system?

OpenStudy (anonymous):

\[\lambda=1\]

OpenStudy (anonymous):

so f sub x - f sub y?

OpenStudy (tkhunny):

Forget that forget that. Setups vary. I was thinking upside down. It's find the way it is.

OpenStudy (anonymous):

so how do i find the critical points? or were they found above?

OpenStudy (tkhunny):

What do you think? With z dropping out, we do not get a unique solution. The 3rd system leads to \(\lambda = -1\). This gives values for x and y. Where does that leave us?

OpenStudy (anonymous):

lol how do i make lambda w/o clicking on "Equation" below?

OpenStudy (anonymous):

this is as far as i've gotten

OpenStudy (tkhunny):

Square x and y and then we can think about what it means.

OpenStudy (anonymous):

am i squaring \[x=-\frac{ 1 }{ 2 }\] and

OpenStudy (anonymous):

\[y=-1\]

OpenStudy (tkhunny):

That is what your equation #4 suggests.

OpenStudy (anonymous):

OOOOOOHHHHH

OpenStudy (anonymous):

\[z=-\frac{ 1 }{ 4 }\]

OpenStudy (tkhunny):

You have it. We were just a little freaked out when equation #3 lost its z . Don't let that scare you next time. Just go with it! It solves how it solves. I think we're done.

OpenStudy (anonymous):

THANK YOU, WISH I COULD GIVE MORE THAN 1 MEDAL

OpenStudy (tkhunny):

This is a tough section. The setup is so simple, but most examples are TOO simple. Real problems rarely cooperate quite so nicely as the examples.

OpenStudy (anonymous):

yes it was very simple now that you explained it. i should've noticed the square's in equation 4. thank you again